OpenStudy (anonymous):

Derive the equation of the parabola with a focus at (−5, −5) and a directrix of y = 7. f(x) = −one twenty-fourth (x − 1)2 − 5 f(x) = one twenty-fourth (x − 1)2 − 5 f(x) = −one twenty-fourth (x + 5)2 + 1 f(x) = one twenty-fourth (x + 5)2 + 1

OpenStudy (welshfella):

if you take any point (x,y) on the curve the distance between the focus and that point = the perpendicular distance between the point and the directrix (definition of a parabola) first distance = sqrt[ (y - (-5))^2 + (x - (-5))^2] = sqrt [ (x +5)^2 + (y + 5)^2] and second distance is y - 7

OpenStudy (welshfella):

so we have sqrt [ (x +5)^2 + (y + 5)^2] = (y - 7 square both sides (x +5)^2 + (y + 5)^2 = (y - 7)^2 simplify this and you'll get the required equation